# CodeIF: Benchmarking the Instruction-Following Capabilities of Large Language Models for Code Generation

### v2

@@ 5.2 Cross-Language Performance Analysis of LLMs
Figure 8 compares the performance of leading LLMs across four programming languages: C++, Java, Python, and Go, highlighting key trends at both the model and language levels.
At the model level, DeepSeek-V3 consistently outperforms others, achieving the highest CCSR in C++ (0.725), Java (0.753), and Go (0.722), with an RSR of 0.787 in Java, likely benefiting from pretraining optimizations for code continuity. Claude-3-5-Sonnet excels in Java, attaining the highest CSR (0.504) and RSR (0.749), indicating specialized mechanisms for handling complex constraints. However, its SSR in Python (0.703) is relatively lower, suggesting challenges in cross-language generalization. GPT-4O demonstrates balanced performance across languages, ranking second in CSR (0.355) and RSR (0.682) for Python, with minimal variation across different programming paradigms (coefficient of variation CV = 0.18).
At the language level, C++ presents the most significant challenges, likely due to its complex template metaprogramming, which increases constraint difficulty. Java exhibits the highest inter-model variance, with Claude-3-5-Sonnet standing out, possibly due to optimizations tailored to its strong type system. Go achieves the highest overall SSR, benefiting from its simple syntax, but RSR fluctuates significantly, reflecting differences in how models handle concurrency constraints.
These findings highlight variations in LLMs’ ability to generalize across languages and suggest potential areas for optimization, such as improving handling of dependency constraints and enhancing consistency across different programming paradigms.
@@ 5.3 Analysis of Instruction Adherence Deviations
Analysis of model-generated responses reveals frequent deviations from specific instructions, particularly in naming conventions and formatting constraints. Models often overlook global formatting rules, such as line limits and character restrictions, when handling general structure instructions.
Moreover, adherence to naming conventions remains inconsistent. For instance, even when explicitly instructed to use PascalCase for variable names, models frequently generate results in lowercase or underscore-separated formats (e.g., transforming current_power into CurrentPower incorrectly).
A particularly notable issue is the tendency to disregard prohibitive instructions. For example, when explicitly instructed to avoid if statements in decision-making—favoring ternary operators or data structures (e.g., dictionaries)—models still generate outputs containing if statements, indicating gaps in constraint enforcement.
5.4 Enhance Model Instruction Compliance
Insights from Appendix Table 5 highlight key strategies for improving instruction adherence in code generation. Supervised Fine-Tuning (SFT) has proven effective, particularly in the Llama series, while model scale remains a crucial factor—larger models like Qwen2.5-72B-Instruct consistently outperform smaller counterparts such as Qwen2.5-3B-Instruct in instruction-following accuracy.
To enhance adherence, several approaches can be employed. First, classifying instructions by priority ensures strict enforcement of hard constraints (e.g., syntax restrictions) while allowing flexibility for soft guidelines (e.g., coding styles). Patterned code generation can provide structured alternatives, such as replacing conditional statements with lookup tables or state machines. Integrating a naming convention engine allows automated conversion of variable and class names (e.g., transforming snake_case to PascalCase). Additionally, leveraging Abstract Syntax Tree (AST) analysis helps detect and transform prohibited structures, such as replacing for loops with while loops or recursive calls. Finally, conflict resolution mechanisms can be introduced to handle contradictory user instructions, such as providing alternative solutions when certain language features are unavailable (e.g., requesting switch-case in Python).
@@ 6 Related Work
Code generation and instruction-following are pivotal capacities under examination in AI research (Feng et al., 2020; Sun et al., 2024; Luo et al., 2024; Wang et al., 2023b; Kim et al., 2018; Li et al., 2023; Lu et al., 2021; Li et al., 2022; Wei et al., 2023; Nijkamp et al., 2023b; Zhuo et al., 2024; Jain et al., 2024; Nijkamp et al., 2023a; Zhang et al., 2023; Allal et al., 2023; Lozhkov et al., 2024a; Roziere et al., 2023; Lozhkov et al., 2024b; Wang et al., 2021; Yan et al., 2023). Several benchmarks have been devised to appraise these capabilities in large-scale models. For code generation, benchmarks like McEval (Chai et al., 2024), FullStackBench (Liu et al., 2024), Repocoder (Zhang et al., 2023), Repobench (Liu et al., 2023), and LiveCodeBench (Jain et al., 2024) have been notable. Similarly, instruction-following capacities are gauged through benchmarks such as InfoBench (Qin et al., 2024), CFBench (Zhang et al., 2024), Instruct-following (Zhou et al., 2023), and FollowBench (Jiang et al., 2024), each tailored to assess different aspects of following instructions given to models.

### v3

@@ 5.2 Cross-Language Performance Analysis of LLMs
Figure 6 compares the performance of leading LLMs across C++, Java, Python, and Go, highlighting trends at both model and language levels. At the model level, DeepSeek-V3 leads with the highest CCSR in C++ (0.725), Java (0.753), and Go (0.722), and an RSR of 0.787 in Java. Claude-3-5-Sonnet excels in Java with the highest CSR (0.504) and RSR (0.749), but shows lower SSR in Python (0.703). GPT-4O demonstrates balanced performance, ranking second in Python’s CSR (0.355) and RSR (0.682), with minimal variance (CV = 0.18). At the language level, C++ is the most challenging due to complex template metaprogramming. Java shows high inter-model variance, with Claude-3-5-Sonnet performing best. Go achieves the highest SSR but fluctuates in RSR. These results highlight cross-language generalization differences and suggest optimization areas like dependency handling and paradigm consistency.
@@ 5.3 Analysis of Instruction Adherence Deviations
Analysis of model-generated responses shows frequent deviations from instructions, especially in naming conventions and formatting constraints. Models often ignore global formatting rules, such as line limits, and inconsistently follow naming conventions. For example, when instructed to use PascalCase, models sometimes output lowercase or underscore-separated formats (e.g., incorrectly transforming current_power into CurrentPower). A notable issue is the disregard for prohibitive instructions. For instance, models often use if statements despite being instructed to avoid them in favor of ternary operators or data structures like dictionaries, revealing gaps in constraint enforcement.
5.4 Improving Instruction Compliance
Appendix Table 5 highlights strategies to improve adherence. Supervised Fine-Tuning (SFT) proves effective, especially in the Llama series, while larger models like Qwen2.5-72B-Instruct outperform smaller ones in instruction-following accuracy. Key improvements include prioritizing hard constraints (e.g., syntax rules) over soft guidelines (e.g., coding styles). Patterned code generation can replace conditional statements with lookup tables or state machines. A naming convention engine can automate variable name formatting (e.g., converting snake_case to PascalCase). Abstract Syntax Tree (AST) analysis can detect and transform prohibited structures, such as replacing for loops with while loops. Conflict resolution mechanisms can address contradictory instructions, offering alternative solutions when certain language features are unavailable (e.g., using Python’s alternatives to switch-case).
@@ 6 Related Work
Code generation and instruction-following are pivotal capacities under examination in AI research Feng et al. (2020); Sun et al. (2024); Luo et al. (2024); Wang et al. (2023b); Kim et al. (2018); Li et al. (2023); Lu et al. (2021); Li et al. (2022); Wei et al. (2023); Nijkamp et al. (2023b); Zhuo et al. (2024); Jain et al. (2024); Nijkamp et al. (2023a); Zhang et al. (2023); Allal et al. (2023); Lozhkov et al. (2024a); Roziere et al. (2023); Lozhkov et al. (2024b); Wang et al. (2021); Yan et al. (2023). Several benchmarks have been devised to appraise these capabilities in large-scale models. For code generation, benchmarks like McEval Chai et al. (2024), FullStackBench Liu et al. (2024), Repocoder Zhang et al. (2023), Repobench Liu et al. (2023), and LiveCodeBench Jain et al. (2024) have been notable. Similarly, instruction-following capacities are gauged through benchmarks such as InfoBench Qin et al. (2024), CFBench Zhang et al. (2024), Instruct-following Zhou et al. (2023), and FollowBench Jiang et al. (2024), each tailored to assess different aspects of following instructions given to models.
@@ 7 Conclusion
8 Limitations
Limited Language Support. CodeIF includes key languages like Java, Python, Go, and C++, but excludes popular ones like JavaScript, Ruby, and Swift. Expanding language coverage would improve its applicability in diverse contexts.
Static Evaluation Focus. CodeIF focuses mainly on static code properties, such as structure and naming conventions, while overlooking dynamic factors like runtime behavior, performance, and debugging. Including dynamic evaluation would better reflect real-world development challenges.
Uniform Metric Weighting. The metrics (CSR, SSR, RSR, CCSR) treat all constraints equally, which may not align with practical priorities. For example, syntactic correctness is often more critical than naming conventions. Introducing weighted scoring could enhance the interpretability of model performance.
Appendix A Prompt Template
Prompt for Instruction Generation You are an instruction compliance evaluator, required to assess the instruction compliance ability of large models. Therefore, you need to generate a series of data for the code generation instruction detection of large models. [Input Format]
I will input a series of data, and you need to generate a dictionary based on these data, which includes two fields “question” and “instruction_list”
Original question: {Original question} Original instruction list:
{instruction_list}
Input Explanation
The original question is the original question. It contains the original code generation problem.
The original instruction list is the original instruction list. It contains randomly generated code compliance instructions. Some instructions will contain directive keywords that need to be replaced and are wrapped in {{}}. Return Format
Return a json data, do not have extra output. The returned dictionary contains two fields: “question” and “instruction_list”
The format is as follows:
    "question": "Optimized question",
    "instruction_list": [
            "instruction_id": "id1",
            "instruction": "Instruction 1"
Explanation
“question”: It is the optimized question, which does not contain any directive instructions, only contains the explanation of the original question. It does not contain any restrictions on the code. Move the instructions in the question to the instruction list
“instruction_list”: It is the optimized instruction list. You should optimize according to the meaning of the question. More in line with the meaning of the question. Instead of directly outputting the original instruction list, note that you should replace all directive keywords that need to be replaced and are wrapped in , and the final output should not contain directive keywords that need to be replaced. Generation Requirements
question: Please generate the optimized question based on the following data, which does not contain any directive instructions, only contains the core content of the original question.
instruction_list: Originated from the input original instruction list. If there are instructions that completely conflict with the meaning of the question or instructions that conflict with each other You should delete as little as possible, you should modify more. Please replace according to the content in the original instruction_list, you should delete as little as possible. Unless it is contradictory instructions, or instructions that cannot be achieved at all, if you only need to generate additional code to meet the requirements, you can replace it.
Prompt for Code Generation As a programming assistant, your task is to generate code snippets based on the user question and instructions given below: Please consider the following points while generating the code snippet: - Make sure you follow the user instructions word to word. If the instruction says to use a specific language or a specific method, use exactly that. - Your output should be a valid code snippet in the programming language indicated in the question or the instructions. - Pay close attention to the syntax and formatting rules of the programming language that you are using. The code should be well-formatted and easy to read. - Make sure that the code snippet you are generating is efficient and is not overly complicated. Output Format: The output should be a valid, well-formatted, and efficient code snippet that adheres to the above question and instructions. Task information User Question: {question} Instructions: {instructions_str} Please generate the code snippet based on the above information:
Prompt for Answer Judgment As a programming assistant, your task is to evaluate whether the generated code strictly follows the instructions given in light of the user’s problem and directives. You need to return a list of the same length as the instructions, containing only ’Yes’ and ’No’, indicating whether the model adhered to each specific instruction. Consider the following when making judgments: - You must strictly follow the user’s instructions. If the instruction requires the use of a specific language or method, you must explicitly check if the code utilizes it. - Your output should be a list of the same length as the instructions, containing only ’Yes’ and ’No’. - Pay close attention to the programming language syntax and formatting rules you are evaluating. The code should be neatly organized and easy to read. - The list you generate should be valid and not overly complex. Task Information User question:
{question}
Instructions:
{instructions_str}
Model-generated response:
{generated_code}
Based on the information provided, determine whether the model has followed the instructions, and return a list of the same length as the instructions, containing only ‘Yes’ and ‘No’. Please note!!! Your output should only contain the list, with no other content. The items in the list should only be ‘Yes’ and ‘No’, with no other words included.
Appendix B More Resluts
ID Type Instruction Format Format Keys
1 global Your entire response should be written in {programming_language}, the use of other programming languages is not allowed. ["programming_language"]
2 global Your code lines should not exceed {characters_num} characters. ["characters_num"]
3 global Your code should use global variables. []
4 global Your code should not use global variables. []
5 global Your function should have at most {parameter_count} parameters. ["parameter_count"]
6 global Your code should not have more than {function_count} functions. ["function_count"]
7 global Your code should not have more than {class_count} classes. ["class_count"]
8 global Your code should not use the {keyword} keyword. ["keyword"]
9 global Your function should not exceed {line_num} lines. ["line_num"]
10 global Your answer in total should not exceed {line_num} lines. ["line_num"]
11 global Your code should use the {keyword} keyword. ["keyword"]
12 structural control Your code should use data structure {data_structure}. ["data_structure"]
13 structural control Your code should not use data structure {data_structure}. ["data_structure"]
14 structural control Your code should use for-loop. []
15 structural control Your code should not use for-loop. []
16 structural control Your code should use while-loop. []
17 structural control Your code should not use while-loop. []
18 structural control Your code should use if statement for decision making. []
19 structural control Your code should not use if statement for decision making. []
20 structural control Your code should use switch statement for decision making. []
21 structural control Your code should not use switch statement for decision making. []
22 variable Your code should define a variable named {variable_name}. ["variable_name"]
23 variable Your code should define an enumeration named {enumeration_name} ["enumeration_name"]
24 variable The variable names in your code should follow the {naming_convention} naming convention ["naming_convention"]
25 variable Variable {variable_name}, type should be {variable_type}. ["variable_name", "variable_type"]
26 variable Variable {variable_name}, should be a global variable. ["variable_name"]
27 variable Variable {variable_name}, should not be a global variable. ["variable_name"]
28 variable Variable {variable_name}, the initial value should be {variable_value}. ["variable_name", "variable_value"]
29 variable Variable {variable_name}, should be a constant. ["variable_name"]
30 variable Variable {variable_name} should not be a constant. ["variable_name"]
31 function Your code should include a function named {function_name}. ["function_name"]
32 function The function names in your code should follow the {naming_convention}. naming convention ["naming_convention"]
33 function Your code should not use any functions from the {disallowed_function_list}. ["disallowed_function_list"]
34 interface Your code should define an interface named {interface_name}. ["interface_name"]
35 interface The interface names in your code should follow the {naming_convention} naming convention. ["naming_convention"]
36 class Your code should define a class named {class_name}. ["class_name"]
37 class The class names in your code should follow the {naming_convention} naming convention. ["naming_convention"]
38 file Your code should be organized in a package named {package_name}. ["package_name"]
39 file Your code should import the following libraries {library_list}. ["library_list"]
40 file Your code should use the function {function_name} from the library {library_name}. ["function_name", "library_name"]
41 file Your code should not use the following libraries {disallowed_library_list}. ["disallowed_library_list"]
42 combination You should initialize an object named {object_name} as an instance of the {class_name} class using {parameters_name_list} for initialization. ["object_name", "class_name", "parameters_name_list"]
43 combination You should define an interface named {interface_name} that includes these methods {method_name_list}. ["interface_name", "method_name_list"]
44 combination Your code should define a class named {class_name} that implements the {interface_name} interface. ["class_name", "interface_name"]
45 combination In your code, the class {class_name} should have these properties {properties_name_list}. ["class_name", "properties_name_list"]
46 combination In your code, the class {class_name} should have these methods {method_name_list}. ["class_name", "method_name_list"]
47 combination The function {function_name} should take {parameter_name_list} as parameters. ["function_name", "parameter_name_list"]
48 combination The function {function_name} should return a {return_type} as its result. ["function_name", "return_type"]
49 combination Your code should be organized in a package named {package_name}, which should contain these classes {class_name_list}. ["package_name", "class_name_list"]
50 combination Your code should be organized in a package named {package_name}, which should contain these functions {function_name_list}. ["package_name", "function_name_list"]
Table 4: Constraint Instruction Table
Models CSR SSR RSR CCSR
Full Easy Hard Full Easy Hard Full Easy Hard Full Easy Hard
Qwen2.5-1.5b-instruct 0.034 0.053 0.015 0.265 0.334 0.197 0.222 0.282 0.162 0.181 0.234 0.128
Qwen2.5-coder-1.5b-instruct 0.058 0.086 0.03 0.358 0.436 0.281 0.301 0.371 0.233 0.251 0.314 0.189
Qwen2.5-3b-instruct 0.078 0.109 0.046 0.415 0.489 0.34 0.357 0.432 0.282 0.299 0.364 0.233
Llama-3.2-3b-instruct 0.101 0.137 0.065 0.396 0.473 0.318 0.344 0.419 0.268 0.305 0.375 0.235
GPT-3.5-turbo 0.102 0.14 0.065 0.41 0.467 0.353 0.362 0.42 0.303 0.314 0.369 0.259
Qwen2.5-coder-3b-instruct 0.097 0.142 0.051 0.445 0.529 0.359 0.383 0.464 0.301 0.33 0.401 0.258
Llama-3.1-8b 0.129 0.178 0.08 0.452 0.551 0.353 0.402 0.497 0.306 0.352 0.44 0.263
Llama-3.1-8b-instruct 0.145 0.187 0.102 0.467 0.544 0.388 0.418 0.493 0.34 0.37 0.444 0.295
Qwen2.5-coder-7b-instruct 0.142 0.198 0.087 0.514 0.59 0.438 0.453 0.533 0.373 0.39 0.463 0.318
Ministral-3b 0.127 0.162 0.092 0.526 0.591 0.46 0.458 0.527 0.39 0.4 0.458 0.342
Phi-3.5-mini-128k-instruct 0.154 0.217 0.09 0.514 0.635 0.391 0.456 0.574 0.337 0.405 0.51 0.299
Ministral-8b 0.161 0.205 0.116 0.552 0.614 0.489 0.486 0.552 0.419 0.431 0.49 0.371
Gemma-2-9b-it 0.171 0.21 0.131 0.573 0.642 0.504 0.513 0.587 0.44 0.445 0.508 0.383
Llama-3.1-70b 0.196 0.232 0.16 0.61 0.664 0.555 0.545 0.607 0.482 0.482 0.533 0.43
Qwen2.5-coder-14b-instruct 0.218 0.276 0.16 0.596 0.667 0.525 0.539 0.614 0.463 0.483 0.55 0.416
Qwen2.5-14b-instruct 0.238 0.279 0.198 0.61 0.676 0.543 0.557 0.628 0.486 0.498 0.565 0.431
Gemini-2.0-flash-exp 0.254 0.29 0.218 0.615 0.648 0.583 0.556 0.593 0.518 0.514 0.547 0.481
Gemma-2-27b-it 0.245 0.3 0.19 0.658 0.709 0.607 0.596 0.652 0.54 0.533 0.588 0.478
Llama-3.1-70b-instruct 0.265 0.3 0.229 0.675 0.723 0.627 0.612 0.667 0.556 0.559 0.601 0.516
Qwen2.5-32b-instruct 0.294 0.337 0.251 0.68 0.722 0.638 0.621 0.674 0.568 0.56 0.604 0.515
Codestral-2501 0.28 0.339 0.219 0.683 0.748 0.617 0.621 0.691 0.551 0.571 0.633 0.507
Phi-4 0.312 0.361 0.262 0.698 0.735 0.66 0.635 0.681 0.589 0.589 0.631 0.546
GPT-4o-mini 0.292 0.348 0.237 0.731 0.78 0.682 0.665 0.724 0.606 0.609 0.66 0.559
GPT-4o 0.338 0.392 0.283 0.721 0.77 0.671 0.665 0.721 0.609 0.616 0.668 0.563
Gemini-exp-1206 0.357 0.41 0.303 0.744 0.781 0.707 0.685 0.734 0.636 0.636 0.675 0.597
Gemini-1.5-pro 0.351 0.383 0.318 0.763 0.794 0.732 0.704 0.744 0.663 0.647 0.679 0.615
GPT-4o-2024-11-20 0.383 0.441 0.325 0.748 0.792 0.702 0.689 0.745 0.633 0.65 0.698 0.602
Deepseek-coder 0.41 0.45 0.37 0.805 0.836 0.773 0.749 0.791 0.707 0.699 0.732 0.666
Deepseek-v3 0.414 0.468 0.359 0.821 0.847 0.794 0.764 0.806 0.723 0.712 0.743 0.68
Models Global Structural Control Variable Interface Function Class File Combination
Llama-3.2-1b-instruct 0.186 0.190 0.206 0.144 0.284 0.260 0.198 0.172
Qwen2.5-1.5b-instruct 0.244 0.236 0.221 0.213 0.355 0.315 0.230 0.213
Qwen2.5-coder-1.5b-instruct 0.328 0.304 0.326 0.293 0.436 0.426 0.351 0.304
Qwen2.5-3b-instruct 0.383 0.346 0.412 0.383 0.468 0.481 0.383 0.366
Llama-3.2-3b-instruct 0.344 0.332 0.393 0.376 0.454 0.447 0.363 0.367
GPT-3.5-turbo 0.388 0.344 0.417 0.375 0.467 0.449 0.378 0.352
Qwen2.5-coder-3b-instruct 0.397 0.367 0.438 0.419 0.511 0.507 0.415 0.403
Llama-3.1-8b 0.410 0.355 0.451 0.424 0.500 0.503 0.413 0.413
Llama-3.1-8b-instruct 0.422 0.373 0.482 0.455 0.524 0.499 0.407 0.437
Qwen2.5-coder-7b-instruct 0.479 0.419 0.497 0.502 0.576 0.571 0.492 0.487
Ministral-3b 0.472 0.403 0.527 0.512 0.618 0.609 0.524 0.535
Phi-3.5-mini-128k-instruct 0.461 0.410 0.512 0.531 0.562 0.574 0.485 0.491
Qwen2.5-7b-instruct 0.484 0.425 0.532 0.548 0.616 0.591 0.520 0.520
Ministral-8b 0.497 0.433 0.541 0.570 0.622 0.631 0.527 0.557
Gemma-2-9b-it 0.541 0.498 0.599 0.510 0.659 0.618 0.543 0.511
Llama-3.1-70b 0.558 0.500 0.652 0.653 0.685 0.671 0.545 0.597
Qwen2.5-coder-14b-instruct 0.541 0.467 0.623 0.669 0.645 0.652 0.547 0.594
Qwen2.5-14b-instruct 0.569 0.526 0.652 0.592 0.649 0.644 0.533 0.559
Gemini-2.0-flash-exp 0.555 0.526 0.653 0.666 0.685 0.669 0.564 0.615
Gemma-2-27b-it 0.621 0.569 0.699 0.640 0.722 0.710 0.607 0.637
Llama-3.1-70b-instruct 0.606 0.546 0.722 0.718 0.744 0.738 0.603 0.680
Qwen2.5-32b-instruct 0.637 0.581 0.713 0.712 0.732 0.742 0.601 0.653
Qwen2.5-72b-instruct 0.633 0.570 0.734 0.711 0.727 0.726 0.645 0.686
Codestral-2501 0.617 0.552 0.723 0.718 0.733 0.746 0.651 0.694
Phi-4 0.633 0.586 0.734 0.739 0.721 0.752 0.677 0.710
Llama-3.3-70b-instruct 0.621 0.634 0.733 0.730 0.759 0.738 0.645 0.695
GPT-4o-mini 0.671 0.663 0.787 0.774 0.784 0.783 0.657 0.710
GPT-4o 0.665 0.651 0.742 0.759 0.743 0.759 0.666 0.716
Qwen2.5-coder-32b-instruct 0.683 0.654 0.776 0.763 0.772 0.758 0.695 0.736
Gemini-exp-1206 0.690 0.677 0.780 0.789 0.798 0.809 0.675 0.727
Gemini-1.5-pro 0.718 0.696 0.814 0.800 0.812 0.815 0.672 0.749
GPT-4o-2024-11-20 0.685 0.666 0.784 0.786 0.779 0.785 0.706 0.755
Claude-3-5-sonnet-20241022 0.677 0.678 0.750 0.736 0.742 0.730 0.640 0.692
Deepseek-coder 0.759 0.714 0.850 0.856 0.846 0.847 0.754 0.813
Deepseek-v3 0.780 0.732 0.866 0.876 0.866 0.873 0.762 0.831
Table 6: The performance of various models on CodeIF for different types of instructions
Models CPP Java Python Go
CCS CS SS RS CCS CS SS RS CCS CS SS RS CCS CS SS RS
Llama-3.2-1b-instruct 0.123 0.023 0.185 0.150 0.190 0.037 0.265 0.221 0.179 0.047 0.262 0.223 0.086 0.022 0.117 0.096
Qwen2.5-1.5b-instruct 0.171 0.023 0.250 0.206 0.191 0.026 0.277 0.228 0.197 0.047 0.298 0.257 0.151 0.040 0.216 0.179
Qwen2.5-coder-1.5b-instruct 0.253 0.068 0.348 0.297 0.259 0.055 0.375 0.308 0.263 0.060 0.380 0.328 0.218 0.049 0.309 0.255
Qwen2.5-3b-instruct 0.251 0.046 0.367 0.302 0.310 0.078 0.419 0.367 0.306 0.092 0.435 0.384 0.327 0.093 0.433 0.365
Llama-3.2-3b-instruct 0.284 0.073 0.377 0.313 0.345 0.121 0.435 0.380 0.321 0.112 0.429 0.383 0.244 0.084 0.304 0.265
GPT-3.5-turbo 0.301 0.085 0.388 0.332 0.367 0.134 0.461 0.409 0.265 0.092 0.371 0.334 0.318 0.088 0.412 0.363
Qwen2.5-coder-3b-instruct 0.339 0.103 0.444 0.380 0.338 0.101 0.453 0.391 0.320 0.091 0.446 0.390 0.323 0.093 0.431 0.363
Llama-3.1-8b 0.319 0.115 0.409 0.354 0.366 0.130 0.477 0.420 0.376 0.152 0.485 0.446 0.330 0.110 0.413 0.363
Llama-3.1-8b-instruct 0.328 0.112 0.432 0.375 0.408 0.173 0.503 0.447 0.393 0.147 0.496 0.455 0.325 0.133 0.406 0.365
Qwen2.5-coder-7b-instruct 0.389 0.147 0.505 0.434 0.375 0.118 0.503 0.444 0.400 0.155 0.531 0.475 0.401 0.154 0.516 0.456
Ministral-3b 0.356 0.107 0.473 0.401 0.410 0.150 0.542 0.476 0.404 0.112 0.538 0.481 0.430 0.138 0.544 0.464
Phi-3.5-mini-128k-instruct 0.354 0.108 0.461 0.388 0.426 0.179 0.532 0.478 0.440 0.180 0.559 0.510 0.380 0.131 0.482 0.422
Qwen2.5-7b-instruct 0.401 0.162 0.514 0.448 0.439 0.152 0.559 0.495 0.397 0.147 0.523 0.471 0.429 0.154 0.541 0.485
Ministral-8b 0.400 0.143 0.518 0.439 0.434 0.158 0.560 0.495 0.410 0.142 0.538 0.481 0.494 0.214 0.599 0.532
Gemma-2-9b-it 0.446 0.200 0.560 0.499 0.446 0.164 0.576 0.518 0.380 0.131 0.510 0.456 0.542 0.204 0.678 0.609
Llama-3.1-70b 0.487 0.201 0.598 0.518 0.507 0.232 0.632 0.572 0.425 0.136 0.579 0.521 0.522 0.226 0.635 0.571
Qwen2.5-coder-14b-instruct 0.464 0.224 0.572 0.514 0.478 0.206 0.592 0.535 0.522 0.216 0.653 0.594 0.454 0.235 0.544 0.490
Qwen2.5-14b-instruct 0.481 0.230 0.590 0.533 0.528 0.265 0.639 0.581 0.472 0.188 0.599 0.550 0.511 0.281 0.603 0.557
Gemini-2.0-flash-exp 0.491 0.259 0.587 0.519 0.575 0.309 0.664 0.604 0.468 0.207 0.584 0.533 0.514 0.233 0.619 0.558
Gemma-2-27b-it 0.529 0.271 0.645 0.579 0.551 0.261 0.676 0.616 0.465 0.179 0.604 0.543 0.611 0.289 0.727 0.665
Llama-3.1-70b-instruct 0.535 0.267 0.653 0.578 0.581 0.276 0.685 0.620 0.555 0.251 0.696 0.639 0.557 0.264 0.655 0.596
Qwen2.5-32b-instruct 0.551 0.314 0.655 0.602 0.589 0.330 0.706 0.638 0.522 0.231 0.665 0.609 0.580 0.311 0.690 0.634
Qwen2.5-72b-instruct 0.543 0.297 0.638 0.574 0.580 0.288 0.701 0.633 0.574 0.284 0.702 0.651 0.573 0.249 0.687 0.610
Codestral-2501 0.562 0.307 0.658 0.595 0.583 0.301 0.694 0.632 0.566 0.249 0.693 0.637 0.569 0.261 0.681 0.611
Phi-4 0.570 0.331 0.663 0.601 0.612 0.328 0.719 0.650 0.587 0.295 0.714 0.660 0.577 0.292 0.679 0.613
Llama-3.3-70b-instruct 0.558 0.300 0.652 0.582 0.621 0.348 0.713 0.644 0.572 0.264 0.709 0.653 0.602 0.317 0.712 0.640
GPT-4o-mini 0.582 0.292 0.684 0.615 0.620 0.299 0.738 0.667 0.586 0.261 0.731 0.674 0.661 0.330 0.775 0.707
GPT-4o 0.600 0.337 0.698 0.639 0.652 0.368 0.748 0.693 0.600 0.312 0.723 0.676 0.603 0.332 0.701 0.636
Qwen2.5-coder-32b-instruct 0.633 0.384 0.717 0.658 0.654 0.401 0.753 0.699 0.621 0.342 0.736 0.688 0.624 0.322 0.731 0.661
Gemini-exp-1206 0.640 0.424 0.726 0.672 0.650 0.360 0.755 0.689 0.590 0.290 0.724 0.674 0.677 0.373 0.777 0.710
Gemini-1.5-pro 0.635 0.370 0.741 0.676 0.674 0.379 0.783 0.720 0.610 0.278 0.758 0.706 0.674 0.395 0.764 0.707
GPT-4o-2024-11-20 0.653 0.374 0.741 0.669 0.683 0.434 0.776 0.716 0.612 0.355 0.724 0.682 0.653 0.358 0.747 0.683
Claude-3-5-sonnet-20241022 0.615 0.425 0.684 0.643 0.720 0.504 0.789 0.749 0.611 0.396 0.703 0.674 0.650 0.444 0.716 0.686
Deepseek-coder 0.709 0.441 0.802 0.735 0.731 0.463 0.819 0.764 0.657 0.336 0.791 0.747 0.702 0.403 0.805 0.744
Deepseek-v3 0.725 0.435 0.831 0.762 0.753 0.497 0.839 0.787 0.651 0.315 0.793 0.744 0.722 0.404 0.822 0.76
Table 7: the evaluation results of different languages on CODEIF. The metrics include Consistent Continuity Satisfaction Rate (CCSR), Complete Satisfaction Rate (CSR), Soft Satisfaction Rate (SSR), and Rigorous Satisfaction Rate (RSR).
Appendix C Baselines
We evaluate over 30 language models spanning both open-source architectures and commercial APIs. The Meta Llama 3 Series Touvron et al. (2023) contains Llama-3.2-1B/3B/8B/70B-Instruct variants and Llama-3.3-70B-Instruct. Qwen2.5 Series Yang et al. (2024) encompasses Qwen2.5-1.5B/3B/7B/14B/32B/72B-Instruct with dedicated code generation models Qwen2.5-Coder-1.5B/3B/7B/14B/32B-Instruct Hui et al. (2024). Mistral Series Jiang et al. (2023) includes Mistral-3B, Mistral-8B, and the code-specialized Codestral-2501.
The evaluation covers Microsoft’s Phi-3.5-Mini-128K-Instruct (3.8B) and Phi-4 Abdin et al. (2024), along with Google’s Gemma-2-9B/27B-It Team (2024b). DeepSeek Series incorporates DeepSeek-Coder Guo et al. (2024) and DeepSeek-V3 DeepSeek-AI (2024). Commercial APIs include OpenAI’s GPT-3.5-Turbo, GPT-4O-Mini, GPT-4O-2024-05-13, and GPT-4O-2024-11-20 Achiam et al. (2023); Google’s Gemini-2.0-Flash-Exp, Gemini-Exp-1206, and Gemini-1.5-Pro Team (2024a); plus Anthropic’s Claude-3.5-Sonnet-20241022.
Appendix D More Data Analysis
Figure 7 shows the proportion of each instruction category. Global constraints dominate (22.77%), followed by Variable constraints (18.17%). This distribution reflects CodeIF’s balanced focus on high-level structural coherence and fine-grained variable precision, ensuring comprehensive evaluation of code generation capabilities. Figure 8 compares instruction distribution across difficulty levels.